Please help me solve this problem.
Find four consecutive even integers whose sum is 340.

Answer by checkley77(12844)

Let x, x+2, x+4, x+6 be the 4 integers.
x+x+2+x+4+x+6=340
4x=340-12
4x=328
x=328/4
x=82 ans. for the first integer.
82+2=84 for the second integer.
82=4=86 for the third integer.
82+6=88 for the last integer.
Proof;
82+84+86+88=340
340=340